Suppose that the mean weight of school children’s bookbags is 17.41 pounds, with a standard deviation of 2.2 pounds. Find the probability that the mean weight of a sample of 30 bookbags will exceed 17 pounds.


1
Expert's answer
2022-05-24T10:23:12-0400

Set the weights of school children's book bags as variable x.

Given: "x\thicksim N(17.41,2.2^2)"

"n(x)=30"

Get: "\mu_{\=x}=\mu_x=17.41," "\sigma_{\=x}^2=\frac{\sigma_x^2}{n}=\frac{2.2^2}{30}=(\frac{2.2}{\sqrt{30}})^2"

"\=x\thicksim N(17.41,(\frac{2.2}{\sqrt{30}})^2)"

"P(\=x>17)=1-P(\=x\le17)\\
=1-P(\frac{\=x -17.41}{\frac{2.2}{\sqrt{30}}}\le\frac {17-17.41}{\frac{2.2}{\sqrt{30}}})\\
=1-P(z\le-1.02)=1-Z(-1.02)\\
=1-[1-Z(1.02)]=Z(1.02)\\
=0.8461"

(using z-score table)